### v2

# Bitangent surfaces and involutions of quartic surfaces

Abstract.
We study the congruence of bitangent lines of an irreducible surface in $\mathbb{P}^{3}$ in arbitrary characteristic, with special attention to quartic surfaces with rational double points and, in particular, Kummer quartic surfaces.
Research of the second author is partially supported by JSPS Grant-in-Aid for Scientific Research (A) No.20H00112.
1. Introduction
Let $\mathbb{G}=G_{1}(\mathbb{P}^{3})$ be the Grassmannian of lines in $\mathbb{P}^{3}$. A surface $F\subset G_{1}(\mathbb{P}^{3})$ is called a congruence of lines. Its cohomology class $[F]$ in $H^{4}(\mathbb{G},\mathbb{Z})$ is determined by two numbers: the order $m$ and the class $n$. The order (resp. class) is the number of lines passing through a general point (resp. contained in a general plane) in $\mathbb{P}^{3}$. We will call the pair $(m,n)$ the bidegree of $F$. The number $m+n$ is equal to the degree of $F$ in the Plücker embedding $\mathbb{G}\hookrightarrow\mathbb{P}^{5}$ (cf. Dolgachev [3], Hudson [9], Jessop [11]).
A surface $X$ in $\mathbb{P}^{3}$ defines a congruence of lines in $\mathbb{P}^{3}$ equal to the closure of the set of lines that touch $X$ at two distinct points. It is classically known as the bitangent surface, and we denote it by ${\mathrm{Bit}}(X)$. For a general $X$, its order and class are also classically known, and we will reproduce the calculation in this paper with the novelty of taking care of the case of characteristic $2$.
For example, it is known that, for a general surface of degree $d$, the bidegree of the surface ${\mathrm{Bit}}(X)$ is equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ [20, Volume 1, p. 281]. In particular, the bidegree of the bitangent surface of a general quartic surface is equal to $(12,28).$ Moreover, the surface is a smooth surface of degree $40$ in $\mathbb{P}^{5}$. It is of general type with $p_{g}=45$ and $K^{2}=360$ [24], [26].
Although, for a general smooth quartic $X$, the surface ${\mathrm{Bit}}(X)$ is irreducible, for special quartic surfaces, even smooth ones, it can be reducible. It is known that, if a quartic surface $X$ is smooth and does not contain lines, then ${\mathrm{Bit}}(X)$ is smooth [1]. However, the surface could be reducible even when it is smooth but contains lines. For example, a general Cayley’s symmetroid quartic surface admits a smooth quartic model with reducible bitangent surface. One of its irreducible components is a Reye congruence of bidegree $(7,3)$ [2], [5, 7.4].
Although admitting some mild singular points in $X$ doesn’t change the bidegree, it may drastically change the surface ${\mathrm{Bit}}(X)$, it even may become reducible. For example, when $X$ is realized as the focal surface of a congruence of lines of bidegree $(2,n)$ (for the focal surface, see [4, §2.1]), the surface ${\mathrm{Bit}}(X)$ is always singular although its singular points are ordinary double points. In the case $n=2$, the congruence of lines is a del Pezzo surface of degree $4$ and the focal surface is the famous $16$-nodal Kummer quartic surface $X$ (Kummer [15]), and the number of irreducible components of ${\mathrm{Bit}}(X)$ is equal 22, six of them are of bidegree $(2,2)$ and 16 of them are of bidegree $(0,1)$. They are planes of lines contained in one of the 16 planes intersecting $X$ along a conic, a trope plane.
[…]
bidegree Number
$p\neq 2$ $(2,2)$ 6
$(0,1)$ 16
$p=2$, $(1,1)$ 3
ordinary $(0,1)$ 4
$p=2$, $(1,1)$ 2
$2$-rank 1 $(0,1)$ 2
$p=2$, $(1,1)$ 1
supersingular $(0,1)$ 1
Table 1. Irreducible components of $\mathrm{Bit}(X)$ for Kummer quartic aurfaces $X$
Observe that $(12,28)$ is always a multiple of the sum of the bidegrees of irreducible components of the bitangent surface in characteristic 2. We do not know any explanation of this fact.
One reason for dropping of the order of ${\mathrm{Bit}}(X)$ in characteristic two is explained by the fact that the discriminant polynomial of a binary form is a square, and another is that irreducible components of bidegree $(1,1)$ appear instead of bidegree $(2,2)$. On the other hand, the class, being equal to the number of bitangent lines to a general plane section $H$ of $X$, also drops. It is equal to $7,4,2$, or $1$ if the Hasse-Witt invariant of $H$ is equal to $3,2,1$, or 00, respectively [19], [22]. We show that, in fact, the last case is not realized if $X$ is not a cone, i.e., a supersingular quartic curve can not be realized as a plane section of a non-conical quartic surface.
On the way, we discuss different kinds of involutions on quartic surfaces and their relationship to Cremona transformations. For example, an irreducible component of the bitangent surface ${\mathrm{Bit}}(X)$ with non-zero order and non-zero class defines a birational involution of $X$ with the pairs of tangency points of bitangent lines as its general orbits.
The plan of the paper is as follows. In Section 2, we will reproduce Salmon’s proof for the formula of the bidegree of the bitangent surface of a general surface $X$ of degree $d$ in characteristic 00 (Theorem 2.1) and show it holds also for smooth surfaces or surfaces with rational double points (Corollary 2.3). In Section 3, we discuss the classical results about the bitangent surface of a Kummer quartic surface in characteristic $p=0$. They extend without change to the case $p\neq 2$. In Section 4, we discuss birational involutions of quartic surfaces. In Section 5, we give a proof (due to G. Kemper) of the fact that the discriminant of a binary form in the case $p=2$ is a square (Proposition 5.1). This reduces the order of congruences of bitangent lines to half. Finally, in the last three sections 6, 7 and 8 we determine the bidegrees of congruences of bitangent lines of Kummer quartics surfaces in characteristic two according to ordinary, $2$-rank 1 and supersingular, respectively (Theorems 6.5, 7.1, 8.1), and show that a supersingular quartic curve cannot be realized as a plane section of a non-conical quartic surface (Theorem 8.3).
Throughout the paper, we assume that the base field $𝕜\Bbbk$ is an algebraically closed field of characteristic $p\geq 0$.
We would like to thank the referee for useful comments.
2. Generalities on the surface of bitangent lines
Let $X=V(F_{d})$ be a normal surface of degree $d$ and let ${\mathrm{Bit}}(X)$ be its bitangent surface, as it is defined in the introduction.
Theorem 2.1.
[…]
$m\leq\frac{1}{2}d(d-1)(d-2)(d-3),\quad n=\frac{1}{2}d(d-2)(d^{2}-9).$
Proof.
We follow Salmon [20, Volume 2, p. 281]). Let $q$ be a general point in $\mathbb{P}^{3}$ and let $\ell=\langle q,q^{\prime}\rangle$ be the line containing $q$ and tangent to $X$ at some point $q^{\prime}=[x_{0},y_{0},z_{0},w_{0}]$. Without loss of generality, we may assume that $q=[0,0,0,1]$ and
$F=w^{d}+A_{1}w^{d-1}+\cdots+wA_{d-1}+A_{d},$
where $A_{k}$ are homogeneous forms of degree $k$ in $x,y,z$.
Plugging in the parametric equation $[s,t]\mapsto[sv+tv^{\prime}]$, where $[v]=q,[v^{\prime}]=q^{\prime}$, we get
$f:=F(sv+tv^{\prime})=(s+tw_{0})^{d}+\sum_{i=1}^{d}t^{i}A_{i}(x_{0},y_{0},z_{0})(s+tw_{0})^{d-i}.$
By polarizing, we can rewrite it in the form
$f=\sum_{k+m=d}s^{k}t^{m}P_{q^{k}}(q^{\prime}),$
where $P_{v^{k}}(v^{\prime})=P_{v^{\prime m}}(v)$ is the value of $k$-th polar of $F$ with pole at $q$ at the vector $v^{\prime}$ (or, equivalently, the value of the totally polarized symmetric multilinear form defined by $f$ at $(v,\ldots,v,v^{\prime},\ldots,v^{\prime})$. Since $q^{\prime}\in X$, we get $P_{v^{0}}(v^{\prime})=F(v^{\prime})=0$. Moreover, because $\ell$ is tangent to $X$ at $q^{\prime}$, we obtain $P_{v}(v^{\prime})=0$. Thus, we can rewrite
$f=s^{2}g_{d-2}(s,t).$
The line $\ell$ is tangent to $X$ at some other point if and only if the binary form $g_{d-2}$ of degree $d-2$
$g_{d-2}(s,t)=\sum_{k=0}^{d-2}s^{k}t^{d-2-k}P_{q^{k+2}}(q^{\prime})$
has a multiple root.
Recall that the discriminant polynomial $D(a_{0},\ldots,a_{d})$ of a binary form $\sum_{i=0}^{d}a_{i}u^{d-i}v^{i}$ of degree $d$ is a homogenous polynomial of degree $2(d-1)$. It is also a bi-homogeneous polynomial of bidegree $(d(d-1),d(d-1))$ with respect to the action of $\mathbb{G}_{m}^{2}$ via
$(a_{0},\ldots,a_{d})\mapsto(\lambda^{d}a_{0},\lambda^{d-1}\mu a_{1},\ldots,\lambda\mu^{d-1}a_{d-1},\mu^{d}a_{d}).$
Thus, we obtain that the locus of points $q^{\prime}$ such that the line $\langle q,q^{\prime}\rangle$ is tangent to $X$ at two points including $q^{\prime}$ is contained in the intersection of hypersurfaces of degrees $d,d-1$, and $(d-2)(d-3)$. This implies that the expected number of bitangent lines passing through $q$ is equal to $\frac{1}{2}d(d-1)(d-2)(d-3)$.
The class $n$ of ${\mathrm{Bit}}(X)$ is equal to the number of bitangent lines to a general plane section $H$ of $X$. Since $X$ is normal, it is a smooth plane curve of degree $d$. Their number is well-known classically, and it is equal to $\frac{1}{2}d(d-2)(d^{2}-9)$ (see [3, 5.5.1, the formula (5.33)]). ∎
Let us see that the equality for $m$ is the expected order of ${\mathrm{Bit}}(X)$. This means that it holds for a general surface of degree $d$. To see this, we also introduce the congruence ${\mathrm{Flex}}(X)$, the closure of the locus of lines that intersect $X$ at some point with multiplicity $\geq 3$. Following the proof of the previous theorem, we find that the order of ${\mathrm{Flex}}(X)$ is equal to the degree of the complete intersection $X\cap P_{q}(X)\cap P_{q^{2}}(X)$. It is expected to be equal to $d(d-1)(d-2)$. The class of ${\mathrm{Flex}}(X)$ is equal to the number of flex tangents of $H$. It is equal to the number of intersection points of $H$ with its Hessian curve of degree $3(d-2)$. It is expected to be equal to $3d(d-2)$.
Let ${\mathrm{pr}}_{q}:X\to\mathbb{P}^{2}$ be the projection of $X$ to a general plane. The ramification curve ${\mathrm{Rm}}(q)$ of the projection is equal to $X\cap P_{q}(X)$. Its degree is equal to $d(d-1)$ and its arithmetic genus is equal to $1+\frac{1}{2}d(d-1)(2d-5)$. The branch curve $B(q)$ of ${\mathrm{pr}}_{q}$ is equal to the projection of ${\mathrm{Rm}}(X)$, hence it is of degree $d(d-1)$. If ${\mathrm{Rm}}(q)$ is smooth, then it is isomorphic to the normalization of $B(q)$. A bitangent line containing the center of the projection is a secant line of ${\mathrm{Rm}}(X)$, hence its projection is a singular point of $B(q)$ with at least two different branches. For a quartic surface, it is an ordinary node. A flex line from ${\mathrm{Flex}}(X)$ containing $q$ is a tangent line to $P_{q}(X)$. Its projection is a singular point with at least two equal branches.
We use the following theorem due to Valentine and Viktor Kulikov [14, Theorem 0.1]:
Theorem 2.2.
[…]
Corollary 2.3.
Assume that $X$ is smooth or has only ordinary double points. Then, the bidegrees of ${\mathrm{Bit}}(X)$ and ${\mathrm{Flex}}(X)$ are equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ and $(d(d-1)(d-2),3d(d-2))$, respectively
Proof.
We choose $q$ general enough such that no bitangent or flex line through $q$ passes through a singular point of $X$. The ramification curve ${\mathrm{Rm}}(q)$ of the projection ${\mathrm{pr}}_{q}$ is equal to the intersection $X\cap P_{q}(X)$ of $X$ with its polar with respect to $q$. We know that the jacobian ideal of an ordinary double point $x\in X$ is the maximal ideal $\mathfrak{m}_{X,x}$. This easily implies that $P_{q}(X)=\sum_{i=0}^{3}a_{i}\frac{\partial F_{d}}{\partial x_{i}}$ is smooth $x$ and, under the projection ${\mathrm{pr}}_{q}$, the double points of ${\mathrm{Rm}}(q)$ are mapped bijectively to ordinary double points of $B(q)$ different from the nodes and cusps coming from bitangent and flex lines.
Since ${\mathrm{Rm}}(q)$ is a complete intersection of degrees $(d,d-1)$, its arithmetic genus $p_{a}$ is equal to $p_{a}=1+\frac{1}{2}d(d-1)(2d-5)$. The projection ${\mathrm{pr}}_{q}$ defines a birational isomorphism between the curves ${\mathrm{Rm}}(q)$ and $B(q)$. The geometric genus $g$ of $B(q)$ is equal to $\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa-\delta_{0}$, where $\delta$ is the number of bitangent lines, $\kappa$ is the number of flex lines, and $\delta_{0}$ is the number of nodes of ${\mathrm{Rm}}(q)$. This gives an inequality
$p_{a}=1+\frac{1}{2}d(d-1)(2d-5)$
$=\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\frac{1}{2}d(d-1)(d-2)(d-3)-d(d-1)(d-2)$
$\geq g=\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa.$
Since $\delta\leq\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa\leq d(d-1)(d-2)$, we obtain the equality $\delta=\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa=d(d-1)(d-2)$.
One checks that this number coincides with the sum $\frac{1}{2}d(d-1)(d-2)(d-3)+d(d-1)(d-2)$ of the expected numbers of ordinary nodes and ordinary cusps of $B(q)$, which are the projections of the bitangent and flex line.
In this proof, we silently assumed that the curves ${\mathrm{Rm}}(q)$ and $B(q)$ are irreducible. One can treat the reducible case in a similar manner, by deviding the bitangents and flex lines into subsets corresponding to irreducible components of the curves. We leave it to the reader to finish the proof. ∎
3. Quartic surfaces: $p\neq 2$
In the special case $d=4$, we expect that, for a smooth or nodal quartic surface in characteristic zero, the bidegree $(m,n)$ of ${\mathrm{Bit}}(X)$ is equal to $(12,28)$. Let $S$ be an irreducible component of ${\mathrm{Bit}}(X)$ and let $q_{S}:Z_{S}\to S$ be the restriction of the tautological line subbundle $q:Z\to G_{1}(\mathbb{P}^{3})$ to the congruence $S$. The pre-image of $X$ under the projection $p_{S}:Z_{S}\to\mathbb{P}^{3}$ is a double cover of $S$ defined by the projection $q_{S}$:
[…]
Thus, any irreducible component $S$ of ${\mathrm{Bit}}(X)$ with non-zero order and non-zero class defines a birational involution $\sigma_{S}$ of $X$. Its set of fixed points is equal to the pre-image of the locus of points on $X$ such that there exists a line intersecting $X$ at this point with multiplicity $4$, or contained in it.
Example 3.1.
It follows from Kummer’s classification of congruences of order $2$ without fundamental curves that the moduli space of quartic surfaces with $16\geq\mu\geq 11$ nodes contains an irreducible component such that a general surface from this component has reducible congruence ${\mathrm{Bit}}(X)$. These irreducible components are realized as the focal surfaces of the congruences of order 2 without fundamental curves. The number of irreducible components and their bidegree are given in the following table (see Kummer [15], Sturm [23, Art 348]).
$\mu$ bidegree Number
$16$ $(2,2)$ 6
$(0,1)$ 16
$15$ $(2,3)$ 6
$(0,1)$ 10
$14$ $(2,4)$ 4
$(0,1)$ 6
$(4,6)$ 1
$13$ $(2,5)$ 3
$(0,1)$ 3
$(6,10)$ 1
$12_{\rm I}$ $(2,6)$ 2
$(0,1)$ 1
$(8,15)$ 1
$12_{\rm II}$ $(2,6)$ 3
$(6,10)$ 1
$11$ $(2,7)$ 1
$(10,21)$ 1
Table 2. Irreducible components of $\mathrm{Bit}(X)$
The irreducible components of order 2 are confocal congruences, i.e. congruences of lines with the same focal surface.
The first case is a Kummer quartic surface $X$. It follows that $X$ admits six involutions with quotient isomorphic to a congruence of bidegree $(2,2)$. This is a quartic del Pezzo surface embedded in $G_{1}(\mathbb{P}^{3})$ via its anti-canonical linear system. Each pair from the six congruences intersects at four points. The branch curve has $28=16+12$ ordinary nodes and $24$ cusps. Its geometric genus is equal to $3$. The ramification curve ${\mathrm{Rm}}(q)$ is a curve on $X$ of degree 12 with 16 nodes, its geometric genus is also equal to $3$.
Let us explain how to see these six irreducible components of bidegree $(2,2)$. Recall that $X$ admits a smooth model $\tilde{X}$ as a surface of degree $8$ in $\mathbb{P}^{5}$ (see [3, 10.3.3]). The surface $\tilde{X}$ is a complete intersection of three quadrics in $\mathbb{P}^{5}$. One can choose projective coordinates such that $\tilde{X}$ is given by the following equations:
(3.1) $\sum_{i=1}^{6}z_{i}^{2}=\sum_{i=1}^{6}a_{i}z_{i}^{2}=\sum_{i=1}^{6}a_{i}^{2}z_{i}^{2}=0,$
where $a_{1},\ldots,a_{6}$ are distinct constants. The projective transformations
$[z_{1},z_{2},z_{3},z_{4},z_{5},z_{6}]\to[z_{1},\pm z_{2},\pm z_{3},\pm z_{4},\pm z_{5},\pm z_{6}]$
induce automorphisms of $\tilde{X}$ which generate a $2$-elementary group of order $2^{5}$. It contains a subgroup of index $2$ that consists of the identity and $15$ projective involutions that restrict to symplectic automorphisms of $\tilde{X}$. [^1] This means that the involution leaves invariant a non-zero regular $2$-form on the surface, or, equivalently, the set of its fixed points is non-empty and finite. They are the involutions that change exactly two signs at the coordinates. The quotient surface of $\tilde{X}$ by a symplectic automorphism is a K3 surface with rational double points. The remaining sixteen involutions are decomposed into two sets. One set consists of ten involutions that change exactly three signs at the coordinates. They are fixed-point-free with the orbit space isomorphic to an Enriques surface. Another set consists of six involutions $g_{i}$ that change only one coordinate with the orbit space isomorphic to a del Pezzo surface of degree $4$. These are the involutions we are interested in.
The first quadric $V(\sum z_{i}^{2})$ can be identified with the Grassmannian quadric $\mathbb{G}$ written in Klein coordinates $z_{i}$ corresponding to six apolar line complexes (see, for example, [3, 10.2]). It is known that any automorphism of $G_{1}(\mathbb{P}^{3})$ comes from a projective collineation or correlation. The involutions $g_{i}$ preserve the apolar line complexes $V(z_{i})$, and hence, come from a correlation. They transform points to planes, and therefore, act on $\mathbb{G}$ by transforming the plane of lines through a point (an $\alpha$-plane) to the plane of lines contained in a given plane (a $\beta$-plane).
Lemma 3.2.
Let $g_{i}$ be considered as a birational involution of $X$. Then, the closure of lines spanned by the orbits of $g_{i}$ is an irreducible component of ${\mathrm{Bit}}(X)$.
Proof.
The surface $\tilde{X}$ sits in $G_{1}(\mathbb{P}^{3})$ as a singular surface of the quadratic line complex $Y=V(\sum a_{i}z_{i}^{2})\cap\mathbb{G}$, that is, it parameterizes the lines that are singular points of the intersection of $Y$ with $\alpha$-planes. We can identify each such $\alpha$-plane $\Omega(x)$ with the point $x\in X$. An automorphism $g_{i}$ acts on $X$ by sending $\Omega(x)$ to $\Omega(\Pi)$ Let $x\in X$ be a general point and let $\Pi=\mathbb{T}_{x}(X)$ be the embedded tangent space of $X$ at $x$. As we explained above, $g_{i}$ interchanges two families $\{\Omega(x)\}_{x\in{\bf P}^{3}}$, $\{\Omega(\Pi)\}_{\Pi\subset\mathbb{P}^{3}}$ of planes, and hence,
$g_{i}(\Omega(x))=\Omega(\Pi^{\prime}),\quad g_{i}(\Omega(\Pi))=\Omega(x^{\prime})$
for some $x^{\prime}\in\mathbb{P}^{3}$ and $\Pi^{\prime}\subset\mathbb{P}^{3}$. Since $g_{i}$ acts on $X$ as a birational automorphism, $x^{\prime}\in X$ and $\Pi^{\prime}=\mathbb{T}_{x^{\prime}}(X)$. Since any point of $H_{i}=V(z_{i})$ is fixed by $g_{i}$,
$\emptyset\neq\Omega(x)\cap H_{i}=g_{i}(\Omega(x)\cap H_{i})=\Omega(\Pi^{\prime})\cap H_{i}=\Omega(x)\cap\Omega(\Pi^{\prime}),$
$\emptyset\neq\Omega(\Pi)\cap H_{i}=g_{i}(\Omega(\Pi)\cap H_{i})=\Omega(\Pi^{\prime})\cap H_{i}=\Omega(x^{\prime})\cap\Omega(\Pi).$
These imply that $x\in\Pi^{\prime}$ and $x^{\prime}\in\Pi$. Since the line $\ell=\langle x,x^{\prime}\rangle\subset\Pi$ and $\ell\subset\Pi^{\prime}$, $\ell$ is tangent to $X$ at $x$ and $x^{\prime}$, that is, it is a bitangent line of $X$. Thus, $\ell$ can be identified with the pair $\{x,x^{\prime}=g_{i}(x)\}$, that is, a point of the orbit space $\tilde{X}/(g_{i})$. ∎
The quotient surface $\tilde{X}/(g_{i})$ is given by
$\sum_{j\neq i}(a_{j}-a_{i})z_{j}^{2}=\sum_{j\neq i}(a_{j}^{2}-a_{i}^{2})z_{j}^{2}=0,$
which is a quartic del Pezzo surface $\mathcal{D}$.
Also, note that the fixed locus of the involution $g_{i}$ is a canonical curve of genus $5$ given as the intersection of three diagonal quadrics. It has special group of automorphisms isomorphic to the $2$-elementary group $2^{4}$. It is classically known as a Humbert curve. Its image in the quartic del Pezzo surface $\mathcal{D}$ belongs to $|-2K_{\mathcal{D}}|$. It is the branch curve of the double cover $\tilde{X}\to\mathcal{D}$. It has the distinguished property that any of the 16 lines on $\mathcal{D}$ splits into a pair of lines on $\tilde{X}$.
4. Birational involutions of a quartic surface
Let $\sigma$ be a birational involution of a quartic surface $X$. The closure of lines spanned by orbits of $\sigma$ is an irreducible congruence $S(\sigma)$ of lines. Fix a general point $P\in\mathbb{P}^{3}$, then the order of $S(\sigma)$ is equal to the cardinality of the set $\{x\in\mathbb{P}^{3}:P\in\langle x,\sigma(x)\rangle\}$.
Suppose that $\sigma$ is the restriction of a Cremona involution $T$. Then, the set of points $x\in\mathbb{P}^{3}$ such that $P\in\langle x,T(x)\rangle$ is classically called an isologue of $T$ and $P$ is its center. It is given by the condition that
${\text{rank}}\begin{pmatrix}a&b&c&d\\ \ x_{0}&x_{1}&x_{2}&x_{3}\\ y_{0}&y_{1}&y_{2}&y_{3}\end{pmatrix}<3,$
where $P=[a,b,c,d],x=[x_{0},x_{1},x_{2},x_{3}],T(x)=[y_{0},y_{1},y_{2},y_{3}].$ It is expected to be a curve of degree $12k+$ 1, where $k$ is the algebraic degree of $T$ [10, p. 175]. However, it also includes the locus of fundamental points of $T$, as well as the closure of the locus of fixed points of $T$. The class of $S(\sigma)$ is expected to be equal to #$\#\Pi\cap T(\Pi)\cap X=4k$.
Let $\sigma$ be a birational involution of a quartic surface. The rational map
$\phi_{\sigma}:X/(\sigma)\dasharrow G_{1}(\mathbb{P}^{3}),x\mapsto\langle x,\sigma(x)\rangle$
is of degree $1$ or $2$, since a general line intersects $X$ with multiplicity $4$. There are three possible scenarios:
  • (i)
    $\phi_{\sigma}$ is of degree $2$;
  • (ii)
    $\phi_{\sigma}$ is of degree $1$, a general line $\ell_{x}=\langle x,\sigma(x)\rangle$ intersects $X$ at two points and two fixed points of $\sigma$;
  • (iii)
    $\phi_{\sigma}$ is of degree $1$, a general line $\ell_{x}$ is a bitangent line of $X$.
In the last case, we say that $\sigma$ is a bitangent involution.
Let $S(\sigma)\subset G_{1}(\mathbb{P}^{3})$ be the congruence of lines, the image of the rational map $\phi_{\sigma}$.
Example 4.1.
Here we give an example of an involution of Type (ii). Let $\sigma$ be the restriction of a projective involution $T$ in $\mathbb{P}^{3}$. Suppose that $p\neq 2$ and the fixed locus of $T$ is the union of two skew lines $\ell_{1},\ell_{2}$. It is clear that the lines $\ell_{x}$ are invariant lines of $T$, hence each such a line has two fixed points on it, one on each of the lines $\ell_{1},\ell_{2}$. This shows that the congruence $S(\sigma)$ is of bidegree $(1,1)$ isomorphic to a smooth quadric. Since the set of fixed points $X^{\sigma}$ consists of $8$ isolated fixed points or it is one-dimensional, we obtain that either both lines lie in $X$ or each line intersects $X$ at four distinct points. In the first case, the degree of $\phi_{\sigma}$ is equal to one. Assume $X$ is smooth, then $Y=X/(\sigma)$ is also smooth. The cover $X\to Y=X/(\sigma)$ is ramified over two $(-2)$-curves, and its branch curve is the union of two disjoint smooth rational curves $C_{1},C_{2}$ with self-intersection $-4$. By the adjunction formula, $|-K_{Y}|=\emptyset,|-2K_{Y}|=\{C_{1}+C_{2}\}$. Thus, $Y$ is a Coble surface obtained by blowing up 10 points on a smooth quadric $Q$, the eight intersection points of two nodal quartic curves $\bar{C}_{1},\bar{C_{2}}$ of bidegree $(2,2)$. The pre-images of the ten exceptional curves on $Y$ in $X$ are ten invariant lines lying in $X$ that contain infinitely many orbits of $\sigma$.
Here is a concrete example: let $X=V(F)$, where
$F=x_{3}^{3}x_{0}+x_{3}^{3}x_{1}+x_{3}x_{0}^{3}+x_{3}x_{0}x_{2}^{2}+x_{3}x_{1}^{3}+x_{0}^{3}x_{2}+x_{0}^{2}x_{1}x_{2}+x_{0}x_{2}^{3}+x_{1}^{3}x_{2}+x_{1}x_{2}^{3}$
and $\sigma:[x_{0},x_{1},x_{2},x_{3}]\mapsto[-x_{0},-x_{1},x_{2},x_{3}]$. The set of fixed points of $\sigma$ consists of two skew lines $\ell_{1}=V(x_{0},x_{1})$ and $\ell_{2}=V(x_{2},x_{3})$. Plugging in the parametric equation $[(s-t)u_{0},(s-t)u_{1},(s+t)u_{2},(s+t)u_{3}]$ of a line $\ell_{x}$, where $x=[u_{0},u_{1},u_{2},u_{3}]$ is not a fixed point of $\sigma$, we obtain the expression
$(s^{2}-t^{2})(F(u_{0},u_{1},u_{2},u_{3})s^{2}+2G(u_{0},u_{1},u_{2},u_{3})st+F(u_{0},u_{1},u_{2},u_{3})t^{2}),$
where
$G=-u_{0}^{3}u_{2}-u_{0}^{3}u_{3}-u_{0}^{2}u_{1}u_{2}+u_{0}u_{2}^{3}+u_{0}u_{2}^{2}u_{3}+u_{0}u_{3}^{3}-u_{1}^{3}u_{2}-u_{1}^{3}u_{3}+u_{1}u_{2}^{3}+u_{1}u_{3}^{3}.$
Thus, the union of lines contained in $X$ and intersecting the skew lines $\ell_{1}$ and $\ell_{2}$ is equal to $X\cap V(G^{\prime})$, where $G^{\prime}$ is obtained from $G$ by replacing $u_{i}$ with $x_{i}$.
We check that the surfaces $V(F)$ and $V(G^{\prime})$ intersect transversally at $10$ lines and intersect at the two lines of fixed points with multiplicity $3$. This shows that $Y=X/(\sigma)\to S(\sigma)=Q$ is the blowing down of ten $(-1)$ curves on the Coble surface $Y$.
Example 4.2.
Here is an example of a bitangent involution. Let $W$ be a general web of quadrics in $\mathbb{P}^{3}$ and let $\mathcal{D}(W)\subset W$ be the discriminant quartic of singular quadrics, classically known as a Cayley quartic symmetroid. A Reye line is a line in $\mathbb{P}^{3}$, which is contained in a pencil from $W$. It is known that the set of Reye lines forms an irreducible congruence ${\text{Rey}}(W)$ of bidegree $(7,3)$, a Reye congruence, and each Reye line is bitangent to the Steinerian surface $X$ of $W$, the set of singular points of quadrics from $W$. The surface $X$ admits a fixed-point-free involution $\sigma$ such that $S(\sigma)={\text{Rey}}(W)$. The involution is defined by $x\mapsto\cap_{Q\in W}P_{x}(Q)$, where $P_{x}(Q)$ is the first polar of $Q$ with pole at $x$.
One can also consider the surface of bitangents ${\mathrm{Bit}}(\mathcal{D}(W))$ of the quartic symmentroid $\mathcal{D}(W)$. It follows from [5, Theorem 7.4.7] that it is an irreducible surface with the normalization isomorphic to the Reye congruence ${\text{Rey}}(W)$. So, ${\text{Rey}}(W)$ admits two birational models as congruences of bidegree $(7,3)$ and $(12,28)$.
Note that the degree of the focal surface $\Phi$ of ${\mathrm{Bit}}(X)$ is equal $22m-2+2g-$ 2, where $g$ is the sectional genus of ${\mathrm{Bit}}(X)$. Since $m=12$, the degree $\geq 22$. However, $\Phi$ could be reducible, and $X$ be its irreducible component. So, the existence of a bitangent involution on $X$ implies that the focal surface of ${\mathrm{Bit}}(X)$ is reducible.
Table (2) gives examples of families of quartic surfaces which admit bitangent involutions with one irreducible component of order $2$.
It is clear that, if a quartic surface $X$ admits a bitangent involution $\sigma$, then the congruence of lines $S(\sigma)$ is an irreducible component of ${\mathrm{Bit}}(X)$.
Remark 4.3.
Of course, any birational involution of a quartic surface $X$ defines an irreducible congruence of lines. If its order is equal to one, it lifts to a birational involution of $\mathbb{P}^{3}$. Given a general point $P$ in $\mathbb{P}^{3}$, there is a unique line passing through $P$ and spanned by an orbit of two points $x,\sigma(x)$ on $X$. We define $T(P)$ to be the fourth point on this line such that the pairs $(x,\sigma(x))$ and $(P,T(P))$ are harmonically conjugate.
5. Quartic surfaces: $p=2$
Let us see what could be wrong if we drop the assumption that the characteristic is equal to zero.
Proposition 5.1.
[…]
Proof.
(supplied by G. Kemper) Take a univariate polynomial $f=a_{n}x^{n}+\cdots+a_{1}x+a_{0}$ whose coefficients are indeterminates over the field $\mathbb{F}_{2}$. With $y_{1},\ldots,y_{n}$ further indeterminates, the polynomial $\prod_{i<j}(y_{i}-y_{j})$ is invariant under the symmetric group $\mathfrak{S}_{n}$ (because the characteristic is two). Since the Galois group of $f$ permutes the roots $r_{i}$, this means that $\prod_{i<j}(r_{i}-r_{j})$ lies in $\mathbb{F}_{2}[a_{1},\ldots,a_{n}]$. (In fact, just write $\prod_{i<j}(y_{i}-y_{j})$ in terms of the elementary symmetric polynomials.) So, the discriminant of $f$ is a square. ∎
Corollary 5.2.
The order $m$ of ${\mathrm{Bit}}(X)$ satisfies
$m\leq\frac{1}{4}d(d-1)(d-2)(d-3).$
Definition 5.3.
Let $m(d):=\frac{1}{2}d(d-1)(d-2)(d-3)$ (resp. $m(d):=\frac{1}{4}d(d-1)(d-2)(d-3)$) for $p\neq 2$ (resp. $p=2$). We say that a surface $X$ is a general projection surface if the order of ${\mathrm{Bit}}(X)$ is equal to $m(d)$.
In characteristic zero, we know that any surface with at most rational double points as singularities is a general projection surface. In the next section, we show that it is not true anymore in characteristic $2$.
Example 5.4.
Let $X=V(F)$ be a quartic surface, where
$F=x_{0}^{3}x_{1}+x_{1}^{3}x_{2}+x_{2}^{3}x_{3}+x_{3}^{3}x_{0}.$
It is invariant with respect to the involution
$\sigma:[x_{0},x_{1},x_{2},x_{3}]\mapsto[x_{2},x_{3},x_{0},x_{1}].$
It has one singular point $x_{0}=[1,1,1,1]$ of type $A_{2}$ and four other singular points $[1,\zeta,\zeta^{3},\zeta^{2}]$, where $\zeta^{5}=1,\zeta\neq 1$. The fixed locus $X^{\sigma}$ of $\sigma$ is the line $\ell=V(x_{0}+x_{2},x_{1}+x_{3})$.
For any point $x=[a,b,c,d]\in X$ not lying on $\ell$, the line $\ell_{x}=\langle x,\sigma(x)\rangle$ contains one extra point $[a+c,b+d,a+c,b+d]\in\ell$. It is tangent to $X$ at this point. So, $\sigma$ is an involution of Type (ii).
[…]
$p_{0}=x_{0}+x_{2},\ p_{1}=x_{1}+x_{3},\ p_{2}=x_{0}x_{2},\ p_{3}=x_{1}x_{3},\ p_{4}=x_{0}x_{1}+x_{2}x_{3}.$
They satisfy the relation
$p_{0}^{2}p_{3}+p_{0}p_{1}p_{4}+p_{1}^{2}p_{2}+p_{4}^{2}=0,$
and embed $\mathbb{P}^{3}/(\sigma)$ into $\mathbb{P}(1,1,2,2,2)$ as a weighted homogeneous hypersurface of degree $4$ with the double line $V(p_{0},p_{1},p_{4})$. We can write
$F=p_{4}(p_{3}+p_{0}^{2})+(p_{2}+p_{1}^{2})(p_{0}p_{1}+p_{4}).$
So, the image of $X$ is the intersection of two hypersurfaces of degree $4$ with the trivial canonical sheaf. It is singular along the $V(p_{0},p_{1},p_{4})$. The quotient $Y=X/(\sigma)$ is isomorphic to the normalization of this surface. Note that polynomials $p_{i}$ do not generate the algebra of invariants of the projective coordinate ring of $X$. In fact, $x_{0}(x_{3}^{3}+x_{0}x_{1})$ and $x_{1}(x_{0}^{3}+x_{1}^{2}x_{2})$ are invariant modulo $(F)$. The map $\phi:Y\to S(\sigma)$ is just the projection map given by $(p_{0},p_{1},p_{3})$. It blows down the lines $V(x_{0}+x_{2},x_{1}+x_{3})$ to the point $[0,0,1]$.
6. Kummer quartic surfaces in characteristic $2$: ordinary case
Kummer quartic surfaces in characteristic 2 are divided into three classes according to curves of genus 2 being ordinary, 2-rank 1 or supersingular (e.g. Katsura–Kondō [13]). In this section, we discuss, the simplest case, an ordinary Kummer quartic surface. The Kummer quartic surface $X$ in characteristic $2$ associated with an ordinary genus 2 curve $C$ is given by
(6.1) $X=V(a(x^{2}y^{2}+z^{2}w^{2})+b(x^{2}z^{2}+y^{2}w^{2})+c(x^{2}w^{2}+y^{2}z^{2})+xyzw),$
where $a,b,c$ are non-zero constants. They coincide with the coefficients of the Igusa canonical model of $C$ (Laszlo–Pauly [16], [17], Katsura-Kondō [13]). The Kummer quartic surface $X$ has four singular points
$p_{1}=[1,0,0,0],\ p_{2}=[0,1,0,0],\ p_{3}=[0,0,1,0],\ p_{4}=[0,0,0,1]$
all of which are rational double points of type $D_{4}$ (Shioda [21], Katsura [12]), and $X$ has four tropes defined by the hyperplane sections $x=0,y=0,z=0,w=0$, respectively.
Lemma 6.1.
Let $S$ be the congruence of lines of bidegree $(1,1)$ of rays intersecting the skew lines $\ell_{1}=V(x,y)$ and $\ell_{2}=V(z,w)$ (or $V(x,z)$ and $V(y,w)$, or $V(x,w)$ and $V(y,z)$). Then, each ray of $S$ is a bitangent line of $X$.
Proof.
It is enough to consider the first pair of skew lines. A line passing through a point $q=[x_{0},y_{0},z_{0},w_{0}]$ not on $\ell_{1}$ or $\ell_{2}$ is the intersection of two planes $\langle\ell_{1},q\rangle$ and $\langle\ell_{2},q\rangle$. Its parametric equation is
$[x,y,z,w]=[sx_{0},sy_{0},tz_{0},tw_{0}].$
It intersects $\ell_{1}$ at the point $[0,0,z_{0},w_{0}]$ and intersects $\ell_{2}$ at the point $[x_{0},y_{0},0,0]$. Plugging in these equations into (6.1), we get
(6.2) $a(x_{0}^{2}y_{0}^{2}s^{4}+z_{0}^{2}w_{0}^{2}t^{4})+(b(x_{0}^{2}z_{0}^{2}+y_{0}^{2}w_{0}^{2})+c(x_{0}^{2}w_{0}^{2}+y_{0}^{2}z_{0}^{2})+x_{0}y_{0}z_{0}w_{0})s^{2}t^{2}=0.$
This equation is a square of a quadratic equation, and hence, the assertion holds. ∎
[…]
$[s:t]=[z_{0}w_{0}:x_{0}y_{0}],$
This defines an explicit bitangent involution $\sigma$ of $X$. We see that it is not defined only at singular points $p_{1},\ldots,p_{4}$ of $X$. It is obvious that the involution is the restriction of the Cremona involution
$T_{1}:[x,y,z,w]\dashrightarrow[xzw,yzw,xyz,xyw].$
It is equal to the product of the standard inversion transformation $T$ and the involution
$g_{1}:[x,y,z,w]\mapsto[y,x,w,z].$
Note that $g_{1}$ is induced from a translation of $J(C)$ by a non-zero 2-torsion. There are three non-zero 2-torsions on $J(C)$ and all three bitangent involutions are the products of $T$ and the involutions induced from the translations.
Proposition 6.2.
The surface $X$ admits three bitangent involution $\sigma_{1},\sigma_{2},\sigma_{3}$. The congruence of lines $S(\sigma_{i})$ is equal to the congruence of lines intersecting two skew lines from the previous lemma. The fixed curve $X^{\sigma_{i}}$ is an elliptic curve of degree $4$, and it is cut out set-theoretically by the quadric $V(xy+zw)$.
Proof.
It is enough to consider the first involution defined by the two lines $\ell_{1}=V(x,y)$ and $\ell_{2}=V(z,w)$. We denote it by $\sigma$. We have already proved the first assertion. Its fixed locus of $T_{1}$ is the quadric $Q=V(xy+zw)$. It cuts out $X$ along a complete intersection of two surfaces:
$xy+zw=0,$
$b(x^{2}z^{2}+y^{2}w^{2})+c(x^{2}w^{2}+y^{2}z^{2})+xyzw=0.$
Plugging in $y=zw/x$ in the second equation, we find that the intersection is a quartic curve taken with multiplicity $2$. It passes through the singular points, and these points are nonsingular. We also check (substituting $x=1$) that its projection to the plane is a smooth cubic curve. Thus, $C=X^{\sigma}$ is, set-theoretically, a smooth quartic elliptic curve.
Note that it is analogous to the fact that the fixed locus of any of the six bitangent involutions of a Kummer quartic surface in characteristic different from $2$ is an octic curve cut out by a quartic surface with multiplicity $2$. ∎
Let us now look at the birational map
$\phi_{\sigma}:Y=X/(\sigma)\dasharrow S(\sigma)\cong\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Recall that $\phi_{\sigma}$ is not defined at the singular points of $X$. The pencil of planes $\ell_{1}^{\perp}$ (resp. $\ell_{2}^{\perp}$) with the base line $\ell_{1}$ (resp. $\ell_{2}$) cuts out in $X$ a pencil of plane quartic curves. The involution $\sigma$ acts identically on the parameters of the pencil. Let
$\pi:\tilde{X}\to X$
be the minimal resolution of $X$. The involution $\sigma$ lifts to a biregular involution $\tilde{\sigma}$ of $\tilde{X}$.
Lemma 6.3.
The pencils $\ell_{1}^{\perp}$ and $\ell_{2}^{\perp}$ define two invariant elliptic pencils $|F_{1}|$ and $|F_{2}|$ on $\tilde{X}$. Each pencil has four reducible fibers: two fibers of type $\tilde{D}_{6}$ and two fibers of type $\tilde{A}_{1}^{*}$ (of Kodaira’s type III).
Proof.
We consider only the pencil defined by $\ell_{1}^{\perp}=V(y+tx)$ with parameter $t\in\mathbb{P}^{1}$. We check that, for a general parameter $t$, the plane quartic $S(x,tx,z,w)$ has two cusps at $p_{3}$ and $p_{4}$, hence its geometric genus is equal to $1$. The cusps are the base points of the pencil of quartics. There are special values of the parameter $t$: $0,\infty,a/b,c/b$. For $t=0$ (resp. $t=\infty$), the quartic is a double conic, it passes through the third singular point $p_{1}\in\ell_{2}$ (resp. $p_{2}\in\ell_{2}$). It is one of the four trope-conics on $X$, i.e. conics cut out set-theoretically by one of trope-planes:
$C_{123}$ $=$ $V(axy+bxz+cyz),$
$C_{124}$ $=$ $V(axy+byw+cxw),$
$C_{134}$ $=$ $V(azw+bxz+cxw),$
$C_{234}$ $=$ $V(azw+byw+cyz).$
Let $E_{i}=E_{0}^{(i)}+E_{1}^{(i)}+E_{2}^{(i)}+E_{3}^{(i)},i=1,2,3,4,$ be the exceptional curves over singular points $p_{i}$ of $X$, where $E_{0}^{(i)}$ is the central component. We check that the proper transform of $V(y)$ intersects one of the components $E_{i}^{(j)},i\neq 0,j=1,3,4$. The corresponding fiber of $|F_{1}|$ is of type $\tilde{D}_{6}$. If $t=a/b$, the quartic has one cusp at one of the base points and a cusp followed by infinitely near node at another base points. It defines a fiber of type $\tilde{A}_{1}^{*}$. ∎
Let us look at the orbit space $\tilde{Y}=\tilde{X}/(\tilde{\sigma})$. Recall that each singular point of $X$ is a rational double point of type $D_{4}$ with the exceptional curve $E_{0}^{(i)}+E_{1}^{(i)}+E_{2}^{(i)}+E_{3}^{(i)}$ over $p_{i}$, where $E_{0}^{(i)}$ is the central component. Observe that the transformation $T_{1}$ blows down each conic $C_{ijk}$ to a singular point:
$C_{123}\to p_{3},\ C_{124}\to p_{4},\ C_{134}\to p_{1},\ C_{234}\to p_{2}.$
The involution $\tilde{\sigma}$ interchanges the proper transforms $\tilde{C}_{123}$ of $C_{123}$ with the central component $E_{0}^{(3)}$, and similarly for other trope conics. It acts on the corresponding reducible fiber via the symmetry of its dual graph.
Now, we are ready to describe the birational morphism
$\tilde{\phi}_{\sigma}:\tilde{Y}\to S(\sigma).$
Let $|\mathfrak{f}_{1}|$ and $|\mathfrak{f}_{2}|$ be the rulings of the quadric $Q=S(\sigma)$ corresponding to the family of planes $\ell_{1}^{\perp}$ and $\ell_{2}^{\perp}$. Their pre-images under $\tilde{X}\to\tilde{Y}\to Q$ are the two elliptic pencils from Proposition 6.2. Let $L_{0}^{(1)},L_{\infty}^{(1)}$ (resp. $L_{0}^{(2)},L_{\infty}^{(2)}$) be the lines from $|\mathfrak{f}_{1}|$ (resp. $|\mathfrak{f}_{2}|$) corresponding to the planes $V(x),V(y)\in\ell_{1}^{\perp}$ (resp. $V(z),V(w)\in\ell_{2}^{\perp}$). Their pre-images in $\tilde{X}$ are the reducible fibers of type $\tilde{D}_{6}$ of $|F_{1}|$ (resp. $|F_{2}|)$. The image of the curve $\tilde{X}^{\tilde{\sigma}}$ in $Q$ is a smooth curve $B$ of bidegree $(2,2)$ that passes through the vertices of the quadrangle of lines $L=L_{0}^{(1)}+L_{\infty}^{(1)}+L_{0}^{(2)}+L_{\infty}^{(2)}$.
We leave it to the reader to check the following assertion:
Proposition 6.4.
The morphism $\tilde{\phi}_{\sigma}:\tilde{X}/(\tilde{\sigma})\to S(\sigma)$ is the blow-up of the eight points on $S(\sigma)$: four vertices of the quadrangle of lines $L$ and infinitely near points to them corresponding to the tangent direction of the curve $B$. The pre-images of the two generators in $|\mathfrak{f}_{i}|$ corresponding to the ramification points of the projection map $B\to|\mathfrak{f}_{i}|^{*}\cong\mathbb{P}^{1}$ are reducible fibers of $|F_{i}|$ of type $\tilde{A}_{1}^{*}$.
So far, we have found that the bitangent surface ${\mathrm{Bit}}(X)$ contains three irreducible components of bidegree $(m,n)=(1,1)$. They correspond to three bitangent involutions $\sigma_{i}$. It also contains four irreducible components of bidegree $(0,1)$ corresponding to the trope-conics $C_{ijk}$.
Theorem 6.5.
The surface ${\mathrm{Bit}}(X)$ is of bidegree $(m,n)=(3,7)$ in $G_{1}(\mathbb{P}^{3})$. It consists of $7$ irreducible components listed in above.
Proof.
It is enough to show that the class $n$ of ${\mathrm{Bit}}(X)$ is equal to $3+4=7$. This is the number of bitangent lines contained in a general plane in $\mathbb{P}^{3}$. We have already shown that $n\geq 7$. On the other hand, it is known that the number of bitangent lines of a smooth plane quartic in characteristic $2$ is $7,4,2$ or $1$ if the Hasse-Witt invariant is equal to $3,2,1$ or 00, respectively (Resmini [19], Stöhr and Voloch [22]). The assertion now follows. ∎
Remark 6.6.
Consider the standard inversion transformation
$T:[x,y,z,w]\dashrightarrow[yzw,xzw,xyw,xyz]$
of $\mathbb{P}^{3}$. It restricts to a fixed-point-free involution $\sigma$ of $\tilde{X}$ with $\tilde{X}/(\sigma)$ isomorphic to an Enriques surface. Let us consider the Plücker embedding of the corresponding congruence of lines. The six minors $p_{ij}$ of the matrix
$\left(\begin{array}[]{cccc}x&y&z&w\\ yzw&xzw&xyw&xyz\\ \end{array}\right)$
are given by
$[p_{12},p_{13},p_{14},p_{23},p_{24},p_{34}]$
$=\left[zw(x^{2}+y^{2}),\ yw(x^{2}+z^{2}),\ yz(x^{2}+w^{2}),\ xw(y^{2}+z^{2}),\ xz(y^{2}+w^{2}),\ xy(z^{2}+w^{2})\right],$
which satisfies
$p_{12}p_{34}+p_{13}p_{24}+p_{14}p_{23}=xyzw((x+y)(z+w)+(x+z)(y+w)+(x+w)(y+z))^{2}=0$
and the cubic equation
(6.3) $p_{12}p_{13}p_{23}+p_{12}p_{14}p_{24}+p_{13}p_{14}p_{34}+p_{23}p_{24}p_{34}=0$
(see Emch [8, (33)]).
Lemma 6.7.
Let $p_{0}=[x_{0},y_{0},z_{0},w_{0}]\in X$ be a general point of the Kummer quartic surface. The line $\ell$ passing through $p_{0}$ and $\sigma(p_{0})$ is not a bitangent line of $X$.
Proof.
The line $\ell$ is given by
$x_{0}(y_{0}^{2}+z_{0}^{2})x+y_{0}(x_{0}^{2}+z_{0}^{2})y+z_{0}(x_{0}^{2}+y_{0}^{2})z=x_{0}(y_{0}^{2}+w_{0}^{2})x+y_{0}(x_{0}^{2}+w_{0}^{2})y+w_{0}(x_{0}^{2}+y_{0}^{2})w=0.$
By putting these equations into (6.1), we get
$xy((x_{0}^{2}(y_{0}^{2}+z_{0}^{2})(y_{0}^{2}+w_{0}^{2})/z_{0}w_{0}(x_{0}^{2}+y_{0}^{2})^{2})x^{2}+(y_{0}^{2}(x_{0}^{2}+z_{0}^{2})(x_{0}^{2}+w_{0}^{2})/z_{0}w_{0}(x_{0}^{2}+y_{0}^{2})^{2})y^{2})$
$\equiv xyzw\ ({\rm modulo\ squares}).$
In general, the coefficients of $x^{3}y,xy^{3}$ are not equal to zero, and hence, the assertion holds. ∎
Let $\ell$ be a line passing through $p_{0}$ and $\sigma(p_{0})$. Let $\ell\cap X=\{p_{0},\sigma(p_{0}),p_{1},p_{2}\}$. The image of $\ell$ under $\sigma$ is a cubic curve, and hence, $p_{1},p_{2}$ are not conjugate to $p_{0},\sigma(p_{0})$. Thus, the map $X\dasharrow G_{1}(\mathbb{P}^{3})$ sending $p_{0}$ to $\ell=\langle p_{0},\sigma(p_{0})\rangle$ has degree 2 onto its image. Therefore, the Enriques surface $\tilde{X}/(\sigma)$ can be embedded into $G_{1}(\mathbb{P}^{3})$ satisfying a cubic relation (6.3). This suggests the following question:
Question 6.8.
Is the Enriques surface $\tilde{X}/(\sigma)$ a Reye congruence of bidegree $(7,3)$
7. Kummer quartic surfaces in $p=2$: $2$-rank $1$ case
In this section, we discuss the case of Kummer quartic surfaces associated with curves of $2$-rank 1. The Kummer quartic surface $X_{1}$ associated with a curve of genus two and $2$-rank $1$ is given by
(7.1) $X_{1}=V(\beta^{2}x^{4}+\alpha^{2}x^{2}z^{2}+x^{2}zw+xyz^{2}+y^{2}w^{2}+z^{4}),$
where $\alpha,\beta$ are constants with $\beta\neq 0$ (Duquesne [7, §3]). The surface $X_{1}$ has exactly two singular points
$p_{1}=[0,0,0,1],\ \ p_{2}=[0,1,0,0]$
of type $D_{8}$ (Shioda [21], Katsura [12]) and contains two tropes cutting by the hyperplane section $x=0$ and $z=0$, respectively. Two tropes meet at $p_{1}$ and $p_{2}$.
Consider the skew lines $\ell=V(x,y)$ and $\ell^{\prime}=V(z,w)$. In the same way as in the ordinary case, we see that the lines meeting $\ell,\ell^{\prime}$ are bitangent lines of $X_{1}$. The corresponding bitangent involution of $X_{1}$ is given by the restriction of a Cremona involution
(7.2) $\sigma_{1}:[x,y,z,w]\dashrightarrow[xz^{2},yz^{2},\beta x^{2}z,\beta x^{2}w].$
Thus, ${\mathrm{Bit}}(X_{1})$ contains a smooth quadric surface $S(\sigma_{1})$.
There exists another bitangent involution
(7.3) $\sigma_{2}:[x,y,z,w]\dashrightarrow[x^{2}z,x^{2}w,xz^{2},yz^{2}],$
which is a composite of $\sigma_{1}$ with a projective linear transformation
$\tau:[x,y,z,w]\to[z,w,\beta x,\beta y].$
Note that $\tau$ is induced from the translation by the non-zero 2-torsion of the Jacobian of the curve.
Now, let us consider the Plücker embeddings of the congruence of bitangent lines defined by $\sigma_{2}$. The six minors $p_{ij}$ of the matrix
$\left(\begin{array}[]{cccc}x&y&z&w\\ x^{2}z&x^{2}w&xz^{2}&yz^{2}\\ \end{array}\right)$
are given by
$[p_{12},p_{13},p_{14},p_{23},p_{24},p_{34}]$
[…]
$=[x^{2},0,xz,xz,xw+yz,z^{2}].$
They satisfy
$p_{13}=0,\ p_{14}+p_{23}=0,\ p_{12}p_{34}+p_{14}p_{23}=0.$
Thus, the congruence of lines $S(\sigma_{2})$ is a quadric cone, a special linear section of the Grassmannan quadric.
We now conclude:
Theorem 7.1.
Let $X_{1}$ be the Kummer quartic surface associated with a smooth curve of genus two and of $2$-rank $1$. Then, ${\mathrm{Bit}}(X_{1})$ is of bidegree $(2,4)$ in $G_{1}(\mathbb{P}^{3})$. It consists of two tropes and two quadric surfaces, one is a smooth quadric $S(\sigma_{1})$ and another is a quadric cone $S(\sigma_{2})$.
Proof.
It is enough to show that the class $n=4$, that is, a general hyperplane section of $X_{1}$ is a plane quartic with exactly four bitangent lines. This follows from the result of Stöhr-Voloch [22, p.60]. We can calculate the Hasse-Witt invariant of the hyperplane section which is equal to 2. Then, it follows from [22, p.60] that the number of bitangent lines of the hyperplane section is equal to 4. ∎
Remark 7.2.
Let $\ell_{0}=V(x,z)$ be the line which is the intersection of two trope-hyperplanes. Then, any bitangent line defined by $\sigma_{2}$ meets $\ell_{0}$ as follows. Let $p=[x_{0},y_{0},z_{0},w_{0}]\in\mathbb{P}^{3}$ be a general point and $q=\sigma_{2}(p)$. The bitangent line $\langle p,q\rangle$ is given by
$[x_{0}(s+tx_{0}z_{0}),\ sy_{0}+tx_{0}^{2}w_{0},\ z_{0}(s+tx_{0}z_{0}),\ sw_{0}+ty_{0}z_{0}^{2}]$
where $[s,t]\in\mathbb{P}^{1}$ is a parameter. The line $\langle p,q\rangle$ meets $\ell_{0}$ at $[0,x_{0},0,z_{0}]$ when $[s,t]=[x_{0}z_{0},1]$.
8. Kummer quartic surfaces in $p=2$: supersingular case
In this section, we consider the supersingular case. In this case, the Kummer quartic surface $X_{0}$ is given by the equation
(8.1) $X_{0}=V(x^{3}w+\alpha x^{3}y+x^{2}yz+\alpha^{2}x^{2}z^{2}+xy^{3}+y^{2}w^{2}+z^{4}),$
[…]
Let us consider the following Cremona involution:
(8.2) $T:[x,y,z,w]\dashrightarrow[x^{3},x^{2}y,\alpha x^{3}+x^{2}z+xy^{2},\alpha x^{2}y+x^{2}w+y^{3}].$
It presereves $X_{0}$, and restricts to a bitangent involution $\sigma$ of $X_{0}$. This follows from a direct calculation.
Let us consider the Plücker embeddings of bitangent lines. The six minors $p_{ij}$ of the matrix
$\left(\begin{array}[]{cccc}x&y&z&w\\ x^{3}&x^{2}y&\alpha x^{3}+x^{2}z+xy^{2}&\alpha x^{2}y+x^{2}w+y^{3}\\ \end{array}\right)$
are given by
$[p_{12},p_{13},p_{14},p_{23},p_{24},p_{34}]$
$=\left[0,\ x^{2}(\alpha x^{2}+y^{2}),\ xy(\alpha x^{2}+y^{2}),\ xy(\alpha x^{2}+y^{2}),\ y^{2}(\alpha x^{2}+y^{2}),\ (xw+yz)(\alpha x^{2}+y^{2})\right]$
$=[0,x^{2},xy,xy,y^{2},xw+yz],$
which satisfies
$p_{12}=0,\ p_{14}+p_{23}=0,\ p_{13}p_{24}+p_{14}p_{23}=0.$
Thus, the congruence of lines $S(\sigma)$ is a quadric cone.
Theorem 8.1.
Let $X_{0}$ be the Kummer quartic surface associated with a supersingular curve. Then, ${\mathrm{Bit}}(X_{0})$ is of bidegree $(1,2)$ in $G_{1}(\mathbb{P}^{3})$. It consists of a trope and a quadric cone.
Proof.
It is enough to show that the class $n=2$, that is, a general hyperplane section of $X_{0}$ is a plane quartic with exactly two bitangent lines. This follows from the result of Stöhr-Voloch [22, p.60], too. We can calculate the Hasse-Witt invariant of the hyperplane section which is equal to 1. Then, it follows from [22, p.60] that the number of bitangent lines of the hyperplane section is equal to 2. ∎
Remark 8.2.
Let $\ell_{0}=V(x,y)$. Then, any bitangent line meets $\ell_{0}$ as follows. Let $p=[x_{0},y_{0},z_{0},w_{0}]\in\mathbb{P}^{3}$ be a general point and $q=\sigma(p)$. The bitangent line $\langle p,q\rangle$ is given by
$[x_{0}(s+tx_{0}^{2}),y_{0}(s+tx_{0}^{2}),z_{0}(s+tx_{0}^{2})+tx_{0}(\alpha x_{0}^{2}+y_{0}^{2}),w_{0}(s+tx_{0}^{2})+ty_{0}(\alpha x_{0}^{2}+y_{0}^{2})],$
where $[s,t]\in\mathbb{P}^{1}$ is a parameter. The line $\langle p,q\rangle$ meets $\ell_{0}$ at $[0,0,x_{0},y_{0}]$ when $[s,t]=[x_{0}^{2},1]$.
Theorem 8.3.
Let $X$ be a non-conical normal quartic surface $X$ over a field of characteristic $2$. Then, its general plane section is a plane quartic curve with positive $2$-rank of its Jacobian.
Proof.
A general plane section of $X$ is a smooth quartic curve. Suppose it is supersingular. A plane quartic curve over any field $K$ (not necessary algebraically closed) is isomorphic to a curve
$Q(x,y,z)^{2}+x(y^{3}+zx^{2})=0,$
where $Q$ is a quadratic form with a non-zero coefficient at $z^{2}$ [18, Proposition 2.4]. Its unique bitangent line is $V(x)$.
[…]